# Towards a Generalizable AI for Materials Discovery: Validation through Immersion Coolant Screening

### v1

@@ Abstract
Artificial intelligence (AI) has emerged as a powerful accelerator of materials discovery, yet most existing models remain problem-specific, requiring additional data collection and retraining for each new property. Here we introduce and validate GATE (Geometrically Aligned Transfer Encoder)—a generalizable AI framework that jointly learns 34 physicochemical properties spanning thermal, electrical, mechanical, and optical domains. By aligning these properties within a shared geometric space, GATE captures cross-property correlations that reduce disjoint-property bias—a key factor causing false negatives in multi-criteria screening. To demonstrate its generalizability, GATE—without any problem-specific reconfiguration—was directly applied to the discovery of immersion cooling fluids for data centers, a stringent real-world challenge defined by the Open Compute Project (OCP). Screening billions of candidates, GATE identified 92,861 molecules as promising for practical deployment. Four were experimentally or literarily validated, showing strong agreement with wet-lab measurements and performance comparable to or exceeding a commercial coolant. These results establish GATE as a ready-to-use, generalizable AI platform readily applicable across diverse materials discovery tasks.
@@ Introduction
Building on this limitation, when independently predicted properties are combined to satisfy multiple design criteria, a systematic bias can arise. In this case, compounds that do not truly meet the requirements but are predicted to pass all thresholds (false positives) are more easily retained than compounds whose predictions and actual values both meet the criteria (true positives). We term this disjoint-property bias. It occurs because each property is modeled in isolation, so correlations and trade-offs that naturally emerge from molecular structure are neglected. As the number of criteria increases, the bias becomes more pronounced, yielding candidates that appear promising in silico yet fail experimental validation (Methods and Supplementary Fig. 1).
@@
The selection of these 34 properties was guided by extensive consultation with domain experts across diverse materials development—including chemical, biological, cosmetic, energies, batteries, electronic, and display materials. Each participating group identified physicochemical parameters most critical to their respective design and optimization workflows, which were consolidated to represent a property space with the highest cross-domain utility. Consequently, GATE encompasses thermal, electrical, mechanical, and interfacial properties that collectively underpin a wide spectrum of real-world materials discovery problems. By training on this expert-curated property set, GATE functions as a generalizable AI platform that can be directly deployed across heterogeneous design tasks without retraining or task-specific modification, providing a unified foundation for accelerating diverse materials development processes across domains.
The capability of GATE is demonstrated in the discovery of immersion cooling fluids for data centers, a stringent real-world challenge. Candidate fluids are guided by diverse physicochemical constraints—spanning thermal, mechanical, electrical, and environmental requirements—codified in detail by the Open Compute Project (OCP) guidelines[30]. Using GATE, billions of virtual and purchasable compounds were screened, promising candidates were identified, and their properties were validated through experimental measurements (thermogravimetric analysis, TGA; differential scanning calorimetry, DSC) and literature values.
@@ Results
To apply the OCP guidelines using GATE, we introduced four pragmatic adaptations (Table 1, right): (i) a dielectric constant threshold of $\leq 6$ at 1 kHz, reflecting that dielectric constants decrease with frequency while OCP specifies $\leq 2.3$ in the 20–40 MHz range (see Supplementary Fig. 2); (ii) a flash point threshold of 140 ℃, owing to larger prediction variability; (iii) substitution of melting point for pour point, ensuring fluidity down to 30-30 ℃ since GATE does not provide pour point estimates; and (iv) adaptation of the “Others” criteria—color, toxicity, and odor—to exclude compounds containing halogens (F, Cl, Br, I) or aromatic rings.
@@
To accelerate the discovery of immersion cooling fluids, we established an AI-enabled high-throughput screening workflow (Fig. 2). The process integrates five key stages, from virtual library construction to experimental validation, thereby linking large-scale computational screening with practical materials discovery.
The workflow begins with the construction of a large virtual library, composed of purchasable compounds and their virtually synthesized derivatives generated under mild, single-step reaction conditions (Fig. 2b). Compounds containing undesirable elements or unstable substructures are subsequently removed through structure-based pre-screening (Fig. 2c). AI-based screening is then applied to predict the physicochemical properties of candidate molecules using GATE, benchmarked against the adapted criteria defined in Table 1 (right) (Fig. 2d). Shortlisted compounds then undergo feasibility evaluation, considering commercial availability, cost, and synthetic accessibility (Fig. 2e). Finally, experimental validation confirms the physicochemical properties of selected candidates (Fig. 2f), ensuring their suitability as single-phase immersion cooling fluids.
@@
We analyzed the candidate fluids emerging from the workflow to understand their chemistries, structures, and properties (Fig. 3). As mentioned previously, since the OCP guideline defines dielectric constant $\leq 2.3$ over 20 MHz–40 GHz while GATE predictions are provided at 1 kHz, adopted a pragmatic screening threshold of dielectric constant $\leq 6$ at 1 kHz. This choice is motivated by the well-known decrease of dielectric constant with frequency and is illustrated by literature data for neoprene redrawn in Supplementary Fig. 2. As this adaptation is heuristic rather than a strict derivation, we further partitioned the candidates into sub-criteria to analyze the characteristic features of each group, thereby providing insights into the molecular traits associated with meeting different performance thresholds.
@@
Notably, the screened set includes many compounds structurally related to mineral and synthetic oils as well as natural and synthetic esters, which were highlighted as promising classes in the OCP white paper[36], underscoring the reliability of our AI-driven discovery pipeline. The agreement among GATE predictions (GATE), experimental measurements (Exp.), and literature values (Lit.) across key physicochemical properties further supports the robustness of the approach. Beyond reproducing known classes, the screening also revealed less conventional motifs—identified specifically in the context of immersion cooling fluids for data centers—including phosphorus groups, alcohols, peroxides, nitrogen/boron species, and metal–organic alkoxides. Although the practical viability of these minor classes remains to be validated, their emergence highlights the potential of AI-guided discovery to expand the design space beyond human-established heuristics.
@@
Across the five experimentally or literature-accessible properties, GATE maintained strong predictive accuracy under OOD conditions (Fig. 4b). Pearson correlation coefficients between predictions and reference values ranged from 0.65 (specific heat) to 0.96 (density), exceeding those of STL models for all benchmarked properties. Notably, GATE preserved meaningful correlations even in the most challenging cases—such as decomposition temperature and specific heat, the latter having only 387 reference data points—where STL failed to generalize beyond its training domain (see Fig. 4c).
@@ Discussion
GATE’s generalizability arises from its architecture and training philosophy. The model jointly learns 34 physicochemical properties spanning thermal, electrical, mechanical, and interfacial domains—selected through expert consultation to ensure broad relevance across diverse materials development. This cross-domain learning enables knowledge transfer among correlated properties, allowing the model to predict a wide range of material behaviors without retraining. Building on this foundation, GATE was practically validated through the discovery of immersion cooling fluids for data centers—a stringent, multi-criteria problem defined by the OCP guidelines. Without any model reconfiguration, GATE screened billions of compounds and identified 92,861 candidates meeting these specifications, several of which exhibited experimentally or literarily verified properties comparable to a commercial coolant. Beyond this proof-of-concept, GATE establishes a unified and ready-to-use foundation for accelerating materials discovery across domains, illustrating that its applicability extends far beyond a single case.
@@ Methods
    $P(A^{\uparrow}\cap B^{\downarrow})=P(A^{\uparrow}\cap B^{\downarrow})+\mathrm{Cov}(\mathbf{1}_{A^{\uparrow}},\mathbf{1}_{B^{\downarrow}})\leq P(A^{\uparrow})\,P(B^{\downarrow}),\quad\text{equivalently }\mathrm{Cov}(\mathbf{1}_{A^{\uparrow}},\mathbf{1}_{B^{\downarrow}})\leq 0.$
@@
The surrogate was trained on 0.5% of all possible product combinations (90/10 train/validation split), achieving throughput of 3,000 molecules per second. Importantly, Pearson correlations with GATE predictions exceeded 0.9 across most properties (Supplementary Table 1), demonstrating that the surrogate preserved the predictive patterns of GATE with only minimal loss in accuracy. Because surrogate predictions are inherently less accurate than GATE, we applied them with relaxed thresholds (Supplementary Table 2), ensuring that viable candidates were retained for final evaluation.
@@
The code that supports the findings of this study is available from the corresponding author upon reasonable request.

### v2

@@ Abstract
Artificial intelligence (AI) has emerged as a powerful accelerator of materials discovery, yet most existing models remain problem-specific, requiring additional data collection and retraining for each new property. Here we introduce and validate GATE (Geometrically Aligned Transfer Encoder)—a generalizable AI framework that jointly learns 34 physicochemical properties spanning thermal, electrical, mechanical, and optical domains. By aligning these properties within a shared geometric space, GATE captures cross-property correlations that reduce disjoint-property bias—a key factor causing false positives in multi-criteria screening. To demonstrate its generalizable utility, GATE—without any problem-specific model reconfiguration—applied to the discovery of immersion cooling fluids for data centers, a stringent real-world challenge defined by the Open Compute Project (OCP). Screening billions of candidates, GATE identified 92,861 molecules as promising for practical deployment. Four were experimentally or literarily validated, showing strong agreement with wet-lab measurements and performance comparable to or exceeding a commercial coolant. These results establish GATE as a generalizable AI platform readily applicable across diverse materials discovery tasks.
@@ Introduction
Consequently, when independently predicted properties are combined to satisfy multiple design criteria, a systematic bias can arise. In this case, compounds that do not truly meet the requirements but are predicted to pass all thresholds (false positives) are more easily retained than compounds whose predictions and actual values both meet the criteria (true positives). We term this disjoint-property bias. It occurs because each property is modeled in isolation, so correlations and trade-offs that naturally emerge from molecular structure are neglected. As the number of criteria increases, the bias becomes more pronounced, yielding candidates that appear promising in silico yet fail experimental validation (Methods and Supplementary Fig. 1).
@@
The selection of these 34 properties was guided by extensive consultation with domain experts across diverse materials development—including polymers, electronic materials, batteries, displays, cosmetics, and agricultural chemicals, as well as functional additives. Each participating group identified physicochemical parameters most critical to their respective design and optimization workflows, which were consolidated to define a property space with the highest cross-domain utility. Consequently, GATE encompasses thermal, electrical, mechanical, and interfacial properties that collectively underpin a wide spectrum of real-world materials discovery problems. By training on this expert-curated property set, GATE functions as a generalizable AI platform that can be directly deployed across heterogeneous design tasks without retraining or task-specific modification, providing a unified foundation for accelerating diverse materials development processes across domains.
To validate practicality, we directly applied GATE—without any problem-specific model reconfiguration—to a real-world materials design problem: the discovery of immersion cooling fluids for data centers. Candidate fluids are guided by diverse physicochemical constraints—spanning thermal, mechanical, electrical, and environmental requirements—codified in detail by the Open Compute Project (OCP) guidelines[30]. Using GATE, billions of virtual and purchasable compounds were screened, promising candidates were identified, and their properties were validated through experimental measurements (thermogravimetric analysis, TGA; differential scanning calorimetry, DSC) and literature values.
@@ Results
To apply the OCP guidelines using GATE, we introduced four pragmatic adaptations (Table 1, right): (i) a dielectric constant threshold of $\leq 6$ at 1 kHz, reflecting that dielectric constants decrease with frequency while OCP specifies $\leq 2.3$ in the 20–40 MHz range (see Supplementary Fig. 2); (ii) a flash-point threshold of 140 °C, reflecting higher variance and lower sensitivity of model predictions around the threshold region; (iii) substitution of melting point for pour point, ensuring fluidity down to 30-30 ℃ since GATE does not provide pour point estimates; and (iv) adaptation of the “Others” criteria—color, toxicity, and odor—to exclude compounds containing halogens (F, Cl, Br, I) or aromatic rings.
@@
To accelerate the discovery of immersion cooling fluids, we established an AI-enabled high-throughput screening workflow (Fig. 2). The process integrates five key stages, from virtual library construction to experimental and literature validation, thereby linking large-scale computational screening with practical materials discovery.
The workflow begins with the construction of a large virtual library, composed of purchasable compounds and their virtually synthesized derivatives generated under mild, single-step reaction conditions (Fig. 2b). Compounds containing undesirable elements or unstable substructures are subsequently removed through structure-based pre-screening (Fig. 2c). AI-based screening is then applied to predict the physicochemical properties of candidate molecules using GATE, benchmarked against the adapted criteria defined in Table 1 (right) (Fig. 2d). Shortlisted compounds then undergo feasibility evaluation, considering commercial availability, cost, and synthetic accessibility (Fig. 2e). Finally, experimental and literature validation confirm the physicochemical properties of selected candidates (Fig. 2f), ensuring their suitability as single-phase immersion cooling fluids.
@@
We analyzed the candidate fluids emerging from the workflow to understand their chemistries, structures, and properties (Fig. 3). As mentioned previously, since the OCP guideline defines dielectric constant $\leq 2.3$ over 20 MHz–40 GHz while GATE predictions are provided at 1 kHz, we adopted a pragmatic screening threshold of dielectric constant $\leq 6$ at 1 kHz. This choice is motivated by the well-known decrease of dielectric constant with frequency and is illustrated by literature data for neoprene redrawn in Supplementary Fig. 2. As this adaptation is heuristic rather than a strict derivation, we further partitioned the candidates into sub-criteria to analyze the characteristic features of each group, thereby providing insights into the molecular traits associated with meeting different performance thresholds.
@@
Notably, many screened compounds are structurally related to mineral/synthetic oils and to natural/synthetic esters. These classes were highlighted as promising in the OCP white paper[36], underscoring the reliability of our AI-driven discovery pipeline. The agreement among GATE predictions (GATE), experimental measurements (Exp.), and literature values (Lit.) across key physicochemical properties further supports the robustness of the approach. Beyond reproducing known classes, the screening also revealed less conventional motifs—identified specifically in the context of immersion cooling fluids for data centers—including phosphorus groups, alcohols, peroxides, nitrogen/boron species, and metal–organic alkoxides. Although the practical viability of these minor classes remains to be validated, their emergence highlights the potential of AI-guided discovery to expand the design space beyond human-established heuristics.
@@
Across the five experimentally or literature-accessible properties, GATE maintained strong predictive accuracy under OOD conditions (Fig. 4b). Pearson correlation coefficients between predictions and reference values ranged from 0.65 (specific heat) to 0.96 (density), exceeding those of STL models for all benchmarked properties. Notably, GATE preserved meaningful correlations even in the most challenging cases—such as decomposition temperature and specific heat, the latter having only 387 training data points—where STL failed to generalize beyond its training domain (see Fig. 4c).
@@ Discussion
GATE’s generalizability arises from its architecture and training philosophy. The model jointly learns 34 physicochemical properties encompassing thermal, electrical, mechanical, and interfacial domains. These properties were selected through extensive consultation with domain experts from diverse materials development fields—including polymers, electronic materials, batteries, displays, cosmetics, and agricultural chemicals, as well as functional additives—to capture the key physicochemical parameters most critical to real-world materials design. This cross-domain learning enables knowledge transfer among correlated properties, allowing the model to predict a wide range of material behaviors without retraining. Building on this foundation, GATE was practically validated through the discovery of immersion cooling fluids for data centers—a stringent, multi-criteria problem defined by the OCP guidelines. Without any model reconfiguration, GATE screened billions of compounds and identified 92,861 candidates meeting these specifications, several of which exhibited experimentally or literarily verified properties comparable to a commercial coolant. Beyond this proof-of-concept, GATE establishes a unified and ready-to-use foundation for accelerating materials discovery across domains, illustrating that its applicability extends far beyond a single case.
@@ Methods
    $P(A^{\uparrow}\cap B^{\downarrow})=P(A^{\uparrow})\,P(B^{\downarrow})+\mathrm{Cov}(\mathbf{1}_{A^{\uparrow}},\mathbf{1}_{B^{\downarrow}})\leq P(A^{\uparrow})\,P(B^{\downarrow}),\quad\text{equivalently }\mathrm{Cov}(\mathbf{1}_{A^{\uparrow}},\mathbf{1}_{B^{\downarrow}})\leq 0.$
@@
The surrogate was trained on 0.5% of all possible product combinations (90/10 train/validation split), achieving throughput of 3,000 molecules per second on the same A40 GPU. Importantly, Pearson correlations with GATE predictions exceeded 0.9 across most properties (Supplementary Table 1), demonstrating that the surrogate preserved the predictive patterns of GATE with only minimal loss in accuracy. Because surrogate predictions are inherently less accurate than GATE, we applied them with relaxed thresholds (Supplementary Table 2), ensuring that viable candidates were retained for final evaluation.
@@
The GATE model is deployed as part of the EXAONE Chemical Agent platform developed by LG AI Research. A hosted version of the model is accessible through the EXAONE Showroom (https://showroom.exaone.ai) upon reasonable request for a trial account. The publicly available service may be periodically updated in its architecture and parameters, which could result in minor differences from the predictions reported in this paper.